OpenStudy (anonymous):

A car with a mass of 1.5 × 10³ kilograms is traveling west at a velocity of 22 meters/second. It hits a stationary car with a mass of 9.0 × 10Æ kilograms. If the collision is inelastic, what is the final direction and approximate velocity of the two cars? A. 14 METERS/SECONDS TO THE WEST B. 14 METERS/ SECONDS TO THE EAST C.22 METERS/SECOND TO THE EAST D. 22 METERS/SECONDS TO THE WEST

OpenStudy (stormfire1):

You can deduce the answer here without any calculations based on the choices you have. First, the car is traveling west and hits a stationary car....so its direction at the end will be west since you're not given any addtional information. The problem states it's an inelastic collision so kinetic energy is not conserved but momentum will be. The car will slow down as a result of the collision. The only choice answer in the list that meets these criteria is A.

OpenStudy (anonymous):

THANKS BOO BUT I GOT IT WRONG LOL IM UPLOADING THIS NEXT QUESTION HOLD UP:)

OpenStudy (stormfire1):

That's surprising. What answer did they say was correct? Also, what was the mass of the stationary car in the question (the exponent part is garbled on my screen)?

OpenStudy (stormfire1):

I assume they say the answer was D but the only way for that to be true is if the mass of the second car was 9kg. The formula to get the final velocity is: \[V_{combined}=\frac{M_1V_1+M_2V_2}{M_1+M_2}\]If the second car was only 9kg, you get:\[V_{combined}=\frac{(1.5x10^3kg)(22 m/s)+(9kg)(0)}{1.5*10^3kg + 9kg}=21.87 m/s\]which would be answer D.

OpenStudy (anonymous):

D isnt right either

OpenStudy (anonymous):

m1v1=m2v2 Use this:) not all that other bullpellet. I got the question right with it